A Bibliography of Publications of Friedrich Gebhardt

Friedrich Gebhardt GMD FIT-KI

Sankt Augustin D-53754 Germany

Tel: 02241-14-2660 FAX: 02241-14-2384

E-mail: friedrich.gebhardt@gmd.de (Internet)

11 May 2013 Version 1.10

Abstract

This bibliography records publications of Friedrich Gebhardt.

Title word cross-reference

[p(1 − p)]^{−1/6} [Geb71a]. B(x; 2/3, 2/3) [Geb71a].

14th [BI91].

5th [GV91].

’96 [GS96]. ’97 [KZ97]. 9th [MM90].

Aachen [Boc88]. Abfragesprachen [BGS76]. Ableitung [Geb73b]. Ähnlichkeit [Geb68b]. algorithm [Geb70]. among [Geb89a, Geb91a]. analysis [Boc88, BI91, Did88, Geb71b, Geb88d, Geb88a, Geb89b, Geb90b, Geb90c, Geb91b, GV91, Geb88c]. analyze [Geb88b]. Anfängen [Geb83]. Anforderungen [GBSOW93]. Angaben [Geb73a]. annual [BI91]. Anzeige [GLP75]. applications [TW90]. Applied [GV91]. approximation [Geb71a]. Approximations [Geb66a, Geb69]. Apr [GV91]. Artificial [GS96]. ASMDA [GV91]. assessment [Geb75e, Geb94b, Geb75b]. August [TW90]. Ausreißerproblems [Geb61]. Austria [TW90]. Auswertung [Geb63].

Bank [Geb86]. banks [Geb87a]. Based [BL96, CG94b, CG94a, Geb73a, Geb97c]. basis [Geb75f, Geb76b, Geb95a, Geb97a]. Bayes [Geb61]. Bayeslösungen [Geb61]. beginning [Geb83].

Beiträge [Geb75a, Mül76]. Bemerkungen [FGM^{+}75]. beta [Geb71a]. beta-integral [Geb71a]. between [Geb85]. Bewertung [Geb75b, Geb94b]. binomial [Geb71a]. Bio [Geb63]. Bio-Auswertung [Geb63]. bioassay [Geb63]. Borges [Geb71a]. boundaries [Geb66c]. building [CGSB^{+}96].

Case [BL96, CG94b, CG94a, GL73, Geb95c, Geb95b, Geb96, Geb97c]. Case-Based [BL96, CG94b, CG94a]. Cases [GVGSB97]. Chi [Geb63]. Chi-Quadrat-Schätzfunktionen [Geb63]. chi-square [Geb63]. Choosing [Geb89a, Geb91a]. Classification [Boc88, BI91]. classifications [Geb91c]. clusters [Geb97b]. Codierung [Geb73a]. coding [Geb87b, Geb73a]. Collections [GD83, Geb77]. comparisons [Geb69]. competing [Geb89a, Geb91a]. completely [Geb71b]. Complex [GVGSB97]. COMPSTAT [MM90]. computational [MM90]. computer [Geb73a]. computer-based [Geb73a]. computerunterstütztes [Geb73a]. Computing [Geb75d]. Concurrency [GGV94a, GGV94b]. Conference [Boc88, BI91, TW90]. considerations [Geb89b]. constitutional [Geb75f]. constraints [GGV94a, GGV94b]. Contributions [Geb75a]. control [GGV94a, GGV94b, Geb95a, Geb97a]. convergent [Geb73b]. counterexample [Geb68a]. criteria [GS78]. criterion [Geb94b]. critical [Geb66a]. Critique [GGR73].

Data [Did88, KZ97, Boc88, BI91, Geb73a, GGR73, Geb75f, Geb76b, Geb83, GD83, Geb87a, Geb88b, Geb88d, Geb88c, Geb88a, Geb89b, Geb90b, Geb90c, Geb91b, GV91, BGS76, Geb78, Geb86]. Database [TW90, Geb94a]. Daten [GGR73]. Datenanalyse [Geb88c]. Datenbanken [Geb87a]. Datenbankmaschinen [Geb86]. Datenbasis [Geb75f, Geb76b]. Datenstrukturen [BGS76, Geb78]. Datenverarbeitung [Geb83, GD83]. December [DSE89]. derivative [Geb73b]. Design [GS78, GS96, CGSB^{+}96, CG94a]. Development [BL96, DSE89, FGM^{+}75]. DEXA [TW90]. dialog [Geb75b]. Dialogsystemen [Geb75b]. dimensional [Geb68a]. Discovering [Geb94a]. Discovery [KZ97]. displaying [GLP75]. distributed [Geb64a, Geb66c]. distribution [Geb64a, Geb69, Geb71a, Geb66c]. document [Geb75b, GS78]. documentation [BGS76, Geb73a, GLP75, Geb78, Geb75c, Geb81]. documents [Geb75e]. Dokumentation [Geb75c]. Dokumentationssystem [Geb73a]. Dokumentationssysteme [Geb81, BGS76, Geb78]. Dokumentationssystemen [Mül76, GLP75]. Dokumenten [Geb75b]. Dokumenten-Retrieval [Geb75b]. Dortmund [Müh75]. Dubrovnik [MM90]. Duncan [Geb66a]. Dynamic [Geb95a, Geb97a].

effect [Geb66b]. Eigenschaften [Geb63]. einem [Geb88c]. Einige [Geb63]. election [Geb88b]. Engineering [GGHV95a, GGHV95b, GGHV95b]. Entwicklung [FGM^{+}75]. Erfordernisse [FGM^{+}75]. Ergebnissen [Geb94b]. estimators [Geb63, Geb66b].

European [KZ97]. Evaluation [BL96, CG94b, CG94a]. Expert [DSE89, Geb85, Geb88b, Geb88d, Geb88c, Geb88a, Geb89b, Geb90a, TW90]. Experten [Geb85]. Experten-Systemen [Geb85]. Expertensystem [Geb88c]. exploiting [Geb95b]. explorativen [Geb88c]. Exploratory [Geb90b, Geb88d, Geb88c, Geb88a, Geb89b, Geb90c, Geb91b].

FABEL [GBSOW93, CGSB^{+}96, GBSOW93]. FABEL-Sprache [GBSOW93]. Fachausschuß [Win78]. Fachbereich [Win78]. factor [Geb68b, Geb71b]. factors [Geb71b]. Faktormatrizen [Geb68b]. Fallstudie [GL73]. Federation [Boc88]. Fibonacci [Geb67]. Fibonacci-sequence [Geb67]. Fifth [Did88]. Finding [Geb97b]. First [Boc88, KZ97]. formal [Geb73a]. formalen [Geb73a]. Fourth [BL96, Geb66c]. Fragestellungen [Geb88c]. France [Did88]. frequencies [Geb75f, Geb76b, Geb76a]. Frühjahrstagung [Win78]. function [Geb64a, Geb73b]. funfte [Müh75]. Funktion [Geb73b]. Fuzzy [Geb93a, Geb93a].

gamma [Geb73b, Geb73b]. Gamma-Funktion [Geb73b]. generalizations [Geb89a, Geb91a]. Generating [Geb64a, Geb67]. German [BL96]. Geschichte [GD83]. Gesellschaft [BI91, Deu73]. GI [Müh75, Win78]. Gießen [Win78]. Gliederungseinheiten [GLP75]. GMDS [Win78]. Granada [GV91]. großen [GL73]. größeren [Geb77].

history [GD83]. i.e [Müh75]. IFCS [Boc88]. ihren [GLP75]. Incomplete [Geb71a]. independent [Geb70]. informatics [Did88]. Informatik [Deu73, Win78]. information [FGM^{+}75, GL73, GGR73, Geb75a, Geb85, Geb85]. Information-Retrieval- [Geb85]. Informationssystem [GL73]. Informationssysteme [FGM^{+}75, Geb75a]. Informationssystemen [Hoß76, GGR73]. integral [Geb71a]. Intelligence [GS96]. Interessantheit [Geb94b]. interesting [Geb90a, Geb94a]. Interestingness [Geb94b]. International [Boc88, Did88, GV91, TW90]. inverting [Geb64a].

Jahrestagung [Deu73, Müh75]. June [KZ97]. Juni [Boc88]. JURIS [FGM^{+}75]. juristischen [Mül76]. juristischer [FGM^{+}75, Geb75a, Geb75c]. Juristisches [GL73].

Karlsruhe [Deu73].

Klartextverarbeitung [Win78].

Klassifikation [BI91]. Knowledge [KZ97, BI91, Geb87a, GGHV95a, GGHV95b]. konvergente [Geb73b]. Kriterium [Geb94b]. Kritik [GGR73].

language [Geb95c, GBSOW93, GGV94a, GGV94b]. languages [BGS76, GS78]. large [GL73]. larger [Geb77]. law [Geb75f, Geb76b]. legal [FGM^{+}75, Geb75a, Geb75c, GLP75, GL73]. letter [Geb87b]. Likelihood [Geb63, Geb71b]. Likelihood- [Geb63]. Links [Geb85]. Literaturbericht [Geb87a]. literature [Geb87a]. logarithmic [Geb73b]. logarithmische [Geb73b]. logic [Geb93a]. Logik [Geb93a]. Luxembourg [DSE89].

Machines [Geb86]. Marburg [BI91].

March [BI91]. matrices [Geb68b]. Maximum [Geb63, Geb71b]. mean [Geb66b]. Medizinische [Win78]. Methodical [FGM^{+}75]. Methodik [Geb75a]. Methodische [FGM^{+}75]. methodology [Geb75a]. Methods [Geb95b, Boc88, CG94b, CG94a, CGSB^{+}96]. Minimum [Geb63]. Mining [KZ97]. model [Geb75e, Geb76a]. Modell [Geb76a]. models [Geb90c, Geb91b, GV91]. Möglichkeiten [GLP75]. moments [Geb66c]. MoMo [GGV94a, GGV94b, GGHV95a, GGHV95b, Geb95c]. monotone [Geb70]. multiple [CGSB^{+}96, Geb66a].

non [Geb95a, Geb97a]. non-sequential [Geb95a, Geb97a]. normal [Geb64a, Geb69]. normally [Geb64a, Geb66c]. normalverteilten [Geb66c]. Normen [GLP75]. norms [GLP75]. numbers [Geb64a, Geb67]. numerical [Geb69].

observations [Geb64b]. October [Did88]. Oktober [Müh75]. one [Geb70]. organization [BI91]. oriented [Geb96]. outlier [Geb61]. outlying [Geb64b]. overlapping [Geb91c].

PKDD [KZ97]. Possibilities [GLP75]. practice [Geb93a]. Praxis [Hoß76, Geb93a]. Principles [KZ97]. priorities [Geb95a, Geb97a]. probabilistic [Geb75e]. problem [Geb88c, Geb61]. proc [BI91, GV91]. Proceedings [Did88, KZ97, Boc88, DSE89, MM90, TW90]. processing [Geb83, GD83]. properties [Geb63]. Prospects [Geb88b]. pseudo [Geb67]. pseudo-random [Geb67].

Quadrat [Geb63]. quadruplets [Geb87b]. Querverbindungen [Geb85]. query [BGS76].

random [Geb64a, Geb67]. range [Geb66a]. Raster [Geb93b]. Rasterähnlichkeit [Geb93b]. Realisierung [Hoß76]. Reasoning [BL96, GVGSB97, CG94b]. Rechenzeitbedarf [Geb75d]. references [Geb75c]. regression [Geb70, Geb90c, Geb91b]. Reihe [Geb73b]. related [Boc88]. relevance [Geb75e]. remarks [FGM^{+}75]. requirements [FGM^{+}75, Geb75d, GBSOW93]. results [Geb90a, Geb94b]. retrieval [CG94b, CG94a, Geb75d, GS78, Geb85, Geb95b, Geb96, Geb97c, Geb75d, Geb75b, Geb75b, Geb85]. Retrieval-System [Geb75d]. review [Geb87a]. risk [Geb64b, Geb66b]. rotation [Geb68a].

Sammlungen [GD83]. sample [Geb66c]. samples [Geb66b]. Schätzfunktionen [Geb63]. searching [GLP75]. Selecting [Geb91b, Geb90a]. Selection [Geb91c, Geb90c]. Semantic [Geb87a]. semantics [Geb90b]. Semantisches [Geb87a]. semi [Geb73b]. semi-convergent [Geb73b]. semi-konvergente [Geb73b]. seminar [DSE89]. Sept [MM90]. September [Did88]. sequence [Geb67]. sequential [Geb95a, Geb97a]. series [Geb73b]. several [Geb69]. shuffling [Geb67]. signatures [Geb87b]. significance [Geb66c]. Signifikanzschranken [Geb66c]. similarity [Geb68b, Geb93b]. simple [Geb75e]. site [CGSB^{+}96]. size [GL73]. Sketches [Geb83]. Skizzen [Geb83]. small [Geb66b].

Societies [Boc88]. solution [Geb71b]. solutions [Geb61]. Some [Geb69, Geb89b, Geb64b, Geb66b, Geb71b, Geb63]. spatial [Geb97b]. specifications [GGV94a, GGV94b]. specified [Geb71b]. Sprache [GBSOW93]. Sprachverarbeitung [Mül76]. square [Geb63]. statements [Geb94a]. Statistical [DSE89, Geb88c]. statistics [Geb77, MM90]. Statistische [Geb88c]. Steuerrecht [Geb76b]. Stichprobenmomentes [Geb66c]. stochastic [GV91]. stragglers [Geb66b]. Strategies [Geb88d, Geb64b]. strategy [Geb90a]. Structure [Geb96, Geb95b, Geb97c]. structure-based [Geb97c]. structures [BGS76, Geb78]. studies [Geb95c]. study [GL73]. subdivisions [GLP75]. Suche [GLP75]. superimposed [Geb87b]. supporting [CGSB^{+}96]. Survey [Geb97c]. Symposium [Did88, GV91, KZ97, MM90]. System [BL96, GL73, Geb88c, Geb90a, Geb73a, Geb75d, Geb75d]. Systemen [Geb85]. Systems [DSE89, FGM^{+}75, Geb75b, Geb88b, Geb88d, Geb88a, Geb89b, Geb95a, Geb95b, Geb97a, TW90, BGS76, GGR73, Geb75a, GLP75, Geb78, Geb81, Geb85].

tax [Geb76b]. test [Geb66a]. Text [Geb87b, Geb77]. Texte [Geb73a]. texts [Geb73a]. Textsammlungen [Geb77]. their [GLP75]. Theorie [Geb93a]. theory [Geb93a]. third [Geb66c]. time [Geb75d]. triplets [Geb87b]. Trondheim [KZ97]. two [Geb68a]. two-dimensional [Geb68a]. types [Geb88c]. Typologie [GGR73]. typology [GGR73].

Umfangs [GL73]. use [Geb71a]. usefulness [Geb88a]. using [Geb90b].

V [Did88]. values [Geb66a].

Variablen [Geb66c].

variables [Geb70, Geb66c].

varimax [Geb68a].

varimax-rotation [Geb68a]. Verfassungsrecht [Geb75f]. Versailles [Did88]. Verteilung [Geb66c]. Verweisungen [Geb75c]. Vienna [TW90]. Virtual [CGSB^{+}96].

Wissen [Geb87a].

word [Geb76a, Geb75f, Geb76b, Geb77].

Workshop [BL96].

Worthäufigkeiten [Geb75f, Geb76b, Geb76a]. Wortstatistiken [Geb77].

zum [Geb75b]. zur [FGM^{+}75, Geb75a, GD83, Geb88c, Mül76]. zwischen [Geb85].

References

Barthel:1976:DAF

[BGS76] Thomas Barthel, Friedrich Gebhardt, and Imant Stellmacher. Datenstrukturen und Abfragesprachen für Dokumentationssysteme [English: Data structures and query languages for documentation systems]. In Hoßfeld [Hoß76], pages 89–106. ISBN 3-446-12236-2.

Bock:1991:CDA

[BI91] Hans-Hermann Bock and Peter Ihm, editors. Classification, data analysis, and knowledge organization: proc. 14th annual conference of the Gesellschaft für Klassifikation (Marburg, March 12–14, 1990), Studies in classification, data analysis, and knowledge organization. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1991. ISBN 0-387-53483-0. LCCN QA276.A1 G47 1990.

Burkhard:1996:FGW

[BL96] H.-D. Burkhard and M. Lenz, editors. Fourth German Workshop on Case-Based Reasoning: System Development and Evaluation. Humboldt-Universität, Berlin, Germany, 1996. ISBN ???? LCCN ????

Bock:1988:CRM

[Boc88] Hans H. Bock, editor. Classification and related methods of data analysis: proceedings of the First Conference of the International Federation of Classification Societies (IFCS) (Aachen, Juni 1987). North Holland-Elsevier Science Publishers, Amsterdam, New York, Oxford, 1988. ISBN 0-444-70404-3. LCCN QA278 .I531 1987.

Coulon:1994:ERMb

[CG94a] Carl-Helmut Coulon and Friedrich Gebhardt. Evaluation of retrieval methods in case-based design. Fabel-Report 24, GMD, Sankt Augustin, Germany, September 1994.

Coulon:1994:ERMa

[CG94b] Carl-Helmut Coulon and Friedrich Gebhardt. Evaluation of retrieval methods in case-based reasoning. In Mark Keane, Jean-Paul Haton, and Michel Manago, editors, EWCBR-94: Second European Workshop on Case-Based Reasoning, pages 283–291. AcknoSoft Press, Paris, Chantilly, 1994.

Coulon:1996:VBS

[CGSB^{+}96] Carl-Helmut Coulon, Wolfgang Gräther, Barbara Schmidt-Belz, Angi Voß, Friedrich Gebhardt, Eckehard Groß, and Jörg W. Schaaf. Virtual building site: supporting building design by multiple methods in FABEL. In Gero and Sudweeks [GS96], pages 465–483. ISBN 0-7923-4080-9. LCCN TA174.A7933 1996.

Deussen:1973:GIJ

[Deu73] Peter Deussen, editor. Gesellschaft fur Informatik, 2. Jahrestagung, 1972, Karlsruhe, volume 78 of Lecture Notes in Economics and Mathematical Systems. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1973. ISBN 3-540-06127-4.

Diday:1988:DAI

[Did88] E. Diday, editor. Data analysis and informatics, V: Proceedings of the Fifth International Symposium (Versailles, France, September 29 – October 2, 1987). INRIA, Domaine de Voluceau, Rocquencourt, Le Chesnay, France, 1988.

DSES:1989:DSE

[DSE89] Development of Statistical Expert Systems, proceedings of a seminar, Luxembourg, December 1987, Eurostat News, Special edition. Office for Official Publications of the European Communities, Luxemburg, 1989. LCCN QA276.4 .D48 1989.

Fiedler:1975:MEJ

[FGM^{+}75] Herbert Fiedler, Friedrich Gebhardt, Bernd S. Müller, Joachim Poetsch, Günter Reiner, and Imant Stellmacher. Methodische Erfordernisse juristischer Informationssysteme. Bemerkungen zur Entwicklung von JURIS [English: Methodical requirements for legal information systems: remarks on the development of JURIS]. In Gebhardt [Geb75a], pages 7–98. ISBN 3-8059-0439-8.

Gebhardt:1993:AFR

[GBSOW93] Friedrich Gebhardt, Brigitte Bartsch-Spörl, Wolfgang Oertel, and Jürgen Walther. Anforderungen an die FABEL-Sprache [English: Requirements for the FABEL language]. Fabel-Report 11, GMD, Sankt Augustin, July 1993. 27 pp.

URL ftp://ftp.gmd.de/GMD/
ai-research/Publications/
Fabel/Fabel-11-Sprache.ps.
gz.

Gebhardt:1983:SGD

[GD83] Friedrich Gebhardt and Dolores Dotterweich. Sammlungen zur Geschichte der Datenverarbeitung [English: Collections on the history of data processing]. Arbeitspapiere der GMD 11, GMD, Sankt Augustin, 1983. 26 pp.

Gebhardt:1961:BAB

[Geb61] Friedrich Gebhardt. Bayeslösungen des Ausreißerproblems [English: Bayes solutions to the outlier problem]. PhD thesis, Universität München, München, 1961. 66 pp.

Gebhardt:1963:EEM

[Geb63]  Friedrich Gebhardt.  Einige

Eigenschaften der Maximum Likelihood- und der Minimum Chi-Quadrat-Schätzfunktionen für die Bio-Auswertung [English: Some properties of the maximum likelihood and minimum chi-square estimators for bioassay]. Biometrische Zeitschrift, 5: 252–269, 1963.

Gebhardt:1964:GND

[Geb64a] Friedrich Gebhardt. Generating normally distributed random numbers by inverting the normal distribution function. Math. Comp., 18:302–306, 1964.

Gebhardt:1964:RSS

[Geb64b] Friedrich Gebhardt. On the risk of some strategies for outlying observations. Ann. Math. Stat., 35:1524–1536, 1964.

Gebhardt:1966:ACV

[Geb66a] Friedrich Gebhardt. Approximations to the critical values for Duncan’s multiple range test. Biometrics, 22:179–182, 1966.

Gebhardt:1966:ESR

[Geb66b] Friedrich Gebhardt. On the effect of stragglers on the risk of some mean estimators in small samples. Ann. Math. Stat., 37: 441–450, 1966.

Gebhardt:1966:VSS

[Geb66c] Friedrich Gebhardt. Verteilung und Signifikanzschranken des 3. und 4. Stichprobenmomentes bei normalverteilten Variablen [English: Distribution and significance boundaries of the third and fourth sample moments of normally distributed variables]. Biometrische Zeitschrift, 8:219– 241, 1966.

Gebhardt:1967:GPR

[Geb67] Friedrich Gebhardt. Generating pseudo-random numbers by shuffling a Fibonacci-sequence. Math. Comp., 21:708–709, 1967.

Gebhardt:1968:CTD

[Geb68a] Friedrich Gebhardt. A counterexample to two-dimensional varimax-rotation. Psychometrika, 33:35–36, 1968.

Gebhardt:1968:AF

[Geb68b] Friedrich Gebhardt. Über die Ähnlichkeit von Faktormatrizen [English: On the similarity of factor matrices]. Psychologische Beiträge, 10:591–599, 1968.

Gebhardt:1969:SNC

[Geb69] Friedrich Gebhardt. Some numerical comparisons of several approximations to the normal distribution. J. Amer. Stat. Assoc., 64:1638–1646, 1969.

Gebhardt:1970:AMR

[Geb70] Friedrich Gebhardt. An algorithm for monotone regression with one or more independent variables. Biometrika, 57:263– 271, 1970.

Gebhardt:1971:IBI

[Geb71a] Friedrich Gebhardt. Incomplete beta-integral B(x; 2/3, 2/3) and [p(1 − p)]^{−1/6} for use with Borges’ approximation of the binomial distribution. Journal of the American Statistical Association, 66:189–191, 1971.

Gebhardt:1971:MLS

[Geb71b] Friedrich Gebhardt. Maximum likelihood solution to factor analysis when some factors are completely specified. Psychometrika, 36:155–163, 1971.

Gebhardt:1973:CTF

[Geb73a] Friedrich Gebhardt. Codierung der Texte und formalen Angaben für ein computerunterstütztes Dokumentationssystem [English: Coding texts and formal data for a computer-based documentation system]. Verlag Dokumentation, Pullach, 1973. ISBN 3-7940-3457-0. 94 pp.

Gebhardt:1973:SRL

[Geb73b] Friedrich Gebhardt. Eine semi-konvergente Reihe für die logarithmische Ableitung der Gamma-Funktion [English: A semi-convergent series for the logarithmic derivative of the gamma function]. Zeitschrift für Angewandte Mathematik und Mechanik (ZAMM), 53:354–355, 1973.

Gebhardt:1975:BMJ

[Geb75a] Friedrich Gebhardt, editor.

Beiträge zur Methodik juristischer Informationssysteme [English: Contributions to the methodology of legal information systems]. Number 5 in Beihefte zu Datenverarbeitung im Recht (DVR). Schweitzer, Berlin, 1975. ISBN 3-8059-0439-8. 208 pp.

Gebhardt:1975:BDD

[Geb75b] Friedrich Gebhardt. Bewertung von Dialogsystemen zum Dokumenten-Retrieval [English: Assessment of dialog systems for document retrieval]. In Mühlbacher [Müh75], pages 109– 117. CODEN LNCSD9. ISBN 0-387-07410-4. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA267.A1 L43 no.34. English or German.

Gebhardt:1975:DJV

[Geb75c] Friedrich Gebhardt. Dokumentation juristischer Verweisungen [English: Documentation of legal references]. In Beiträge zur Methodik juristischer Informationssysteme [English: Contributions to the methodology of legal information systems] [Geb75a], pages 157–177. ISBN 3-8059-0439-8.

Gebhardt:1975:RRS

[Geb75d] Friedrich Gebhardt. Rechenzeitbedarf für ein Retrieval-System [English: Computing time requirements for a retrieval system]. In Beiträge zur Methodik juristischer Informationssysteme [English: Contributions to the methodology of legal information systems] [Geb75a], pages 131–141. ISBN 3-8059-0439-8.

Gebhardt:1975:SPM

[Geb75e] Friedrich Gebhardt. A simple probabilistic model for the relevance assessment of documents. Information Processing & Management, 11:59–65, 1975.

Gebhardt:1975:WDV

[Geb75f] Friedrich Gebhardt. Worthäufigkeiten in der Datenbasis Verfassungsrecht [English: Word frequencies in the data basis constitutional law]. In Beiträge zur Methodik juristischer Informationssysteme [English: Contributions to the methodology of legal information systems] [Geb75a], pages 99–130. ISBN 3-8059-0439-8.

Gebhardt:1976:MW

[Geb76a] Friedrich Gebhardt. Ein Modell für Worthäufigkeiten [English: A model for word frequencies]. In Müller [Mül76], pages 35–48. ISBN 3-8059-0457-6.

Gebhardt:1976:WDS

[Geb76b] Friedrich Gebhardt. Worthäufigkeiten in der Datenbasis Steuerrecht [English: Word frequencies in the data basis tax law]. Datenverarbeitung im Recht (DVR), 5: 277–290, 1976.

Gebhardt:1977:WGT

[Geb77]  Friedrich Gebhardt.  Wort-

statistiken an größeren Textsammlungen [English: Word statistics for larger text collections]. Nachrichten für Dokumentation, 28:53–57, 1977.

Gebhardt:1978:DDS

[Geb78] Friedrich Gebhardt. Datenstrukturen für Dokumentationssysteme [English: Data structures for documentation systems]. In Wingert [Win78], pages 74–79. ISBN 3-540-08634-X.

Gebhardt:1981:DSG

[Geb81] Friedrich Gebhardt. Dokumentationssysteme [English: Documentation systems]. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1981. ISBN 3-540-10744-4. 331 pp.

Gebhardt:1983:SAD

[Geb83] Friedrich Gebhardt, editor.

Skizzen aus den Anfängen der Datenverarbeitung [English: Sketches from the beginning of data processing]. Number 143 in Berichte der GMD. Oldenbourg, München, 1983. ISBN 3-486-27971-8. 100 pp.

Gebhardt:1985:QIR

[Geb85]  Friedrich Gebhardt.  Quer-

verbindungen zwischen Information-Retrieval- und Experten-Systemen [English: Links between information retrieval and expert systems]. Nachrichten für Dokumentation, 36:255–263, 1985.

Gebhardt:1986:DBM

[Geb86] Friedrich Gebhardt. Datenbankmaschinen [English: Data Bank Machines]. Number 163 in Berichte der GMD. Oldenbourg, München, 1986. ISBN 3-486-20287-1. 112 pp. LCCN QA1.G34 no.163.

Gebhardt:1987:SWD

[Geb87a] Friedrich Gebhardt. Semantisches Wissen in Datenbanken – ein Literaturbericht [English: Semantic knowledge in data banks: a literature review]. Informatik-Spektrum, 10:79–98, 1987.

Gebhardt:1987:TSS

[Geb87b] Friedrich Gebhardt. Text signatures by superimposed coding of letter triplets and quadruplets. Information Systems, 12: 151–156, 1987.

Gebhardt:1988:UES

[Geb88a] Friedrich Gebhardt. On the usefulness of expert systems in exploratory data analysis. In Diday [Did88], pages 183–188.

Gebhardt:1988:PES

[Geb88b] Friedrich Gebhardt. Prospects for expert systems to analyze election data. In Bock [Boc88], pages 691–696. ISBN 0-444-70404-3. LCCN QA278 .I531 1987.

Gebhardt:1988:SFE

[Geb88c] Friedrich Gebhardt. Statistische Fragestellungen bei einem Expertensystem zur explorativen Datenanalyse [English: Statistical problem types at an expert system for exploratory data analysis]. Number 137 in GMD-Studien. GMD, Sankt Augustin, 1988. ISBN 3-88457-137-0. 123 pp.

Gebhardt:1988:SED

[Geb88d] Friedrich Gebhardt. Strategies for exploratory data analysis in expert systems. In D. Edwards and N. E. Raun, editors, COMPSTAT 88: short communications and posters, pages 81– 82. Physica-Verlag, Heidelberg, Germany, 1988. ISBN 0-387-91336-X (New York), 3-7908-0411-8 (Heidelberg).

Gebhardt:1989:CAC

[Geb89a] Friedrich Gebhardt. Choosing among competing generalizations. Arbeitspapiere der GMD 421, GMD, Sankt Augustin, Germany, 1989. 16 pp.

Gebhardt:1989:SCE

[Geb89b] Friedrich Gebhardt. Some considerations on expert systems in exploratory data analysis. In DSES ’87 [DSE89], pages 303– 315. LCCN QA276.4 .D48 1989.

Gebhardt:1990:ESS

[Geb90a] Friedrich Gebhardt. An expert system strategy for selecting interesting results. In Momirovic and Mildner [MM90], pages 81– 85. ISBN 3-7908-0475-4. LCCN QA276.4 .C55 1990.

Gebhardt:1990:EDA

[Geb90b] Friedrich Gebhardt. Exploratory data analysis using data semantics. In Tjoa and Wagner [TW90], pages 521–526. ISBN 3-211-82234-8, 0-387-82234-8. LCCN QA76.9.E96 T551 1990.

Gebhardt:1990:SRM

[Geb90c] Friedrich Gebhardt. On the selection of regression models in exploratory data analysis. Arbeitspapiere der GMD 466, GMD, Sankt Augustin, Germany, 1990. 16 pp.

Gebhardt:1991:CAC

[Geb91a] Friedrich Gebhardt. Choosing among competing generalizations. Knowledge Acquisition, 3: 361–380, 1991.

Gebhardt:1991:SRM

[Geb91b] Friedrich Gebhardt. Selecting regression models in exploratory data analysis. In Gutiérrez and Valderrama [GV91], pages 262– 273. ISBN 981-02-0644-5. LCCN QA402 .I568 1991.

Gebhardt:1991:SOC

[Geb91c] Friedrich Gebhardt. Selection from overlapping classifications. In Bock and Ihm [BI91], pages 9– 13. ISBN 0-387-53483-0. LCCN QA276.A1 G47 1990.

Gebhardt:1993:FLT

[Geb93a] Friedrich Gebhardt. Fuzzy Logik – Theorie und Praxis [English: Fuzzy logic – theory and practice]. KI – Künstliche Intelligenz, 7(4):71–73, 1993.

Gebhardt:1993:R

[Geb93b] Friedrich Gebhardt. Rasterähnlichkeit [English: Raster similarity]. Fabel-Report 19, GMD, Sankt Augustin, December 1993. 14

pp. URL ftp://ftp.gmd.de/
GMD/ai-research/Publications/
Fabel/Fabel-19-Rasteraehnlichk.
ps.gz.

Gebhardt:1994:DIS

[Geb94a] Friedrich Gebhardt. Discovering interesting statements from a database. Applied Stochastic Models and Data Analysis, 10:1– 14, 1994.

Gebhardt:1994:ICA

[Geb94b] Friedrich Gebhardt. Interessantheit als Kriterium für die Bewertung von Ergebnissen [English: Interestingness as a criterion for the assessment of results]. Informatik – Forschung und Entwicklung, 9:9–21, 1994.

Gebhardt:1995:DPB

[Geb95a]  Friedrich Gebhardt.  Dy-

namic priorities as a basis for control in non-sequential systems. Fabel-Report 37, GMD, Sankt Augustin, Germany, September 1995. 37 pp.

URL ftp://ftp.gmd.de/GMD/
ai-research/Publications/

                                        Fabel/Fabel-37-priorities.
                                        ps.gz.

Gebhardt:1995:MSC

[Geb95b] Friedrich Gebhardt. Methods and systems for case retrieval exploiting the case structure. Fabel-Report 39, GMD, Sankt Augustin, Germany, December 1995. 52 pp.

                                        URL ftp://ftp.gmd.de/GMD/
                                        ai-research/Publications/
                                        Fabel/Fabel-39-case-retr.
                                        ps.gz.

Gebhardt:1995:MLC

[Geb95c] Friedrich Gebhardt. MoMo: language and case studies. Fabel-Report 36, GMD, Sankt Augustin, Germany, July 1995. 59

                                        pp. URL ftp://ftp.gmd.de/
                                        GMD/ai-research/Publications/
                                        Fabel/Fabel-36-MoMo-language.
                                        ps.gz.

Gebhardt:1996:SOC

[Geb96]  Friedrich Gebhardt.  Struc-

ture oriented case retrieval. In Burkhard and Lenz [BL96], pages 95–102. ISBN ???? LCCN ????

Gebhardt:1997:DPB

[Geb97a] Friedrich Gebhardt. Dynamic priorities as a basis for control in non-sequential systems. Cybernetics and Systems, 28(2):121– 161, 1997. CODEN CYSYDH. ISSN 0196-9722.

Gebhardt:1997:FSC

[Geb97b] Friedrich Gebhardt. Finding spatial clusters. In Komorowski and Zytkow [KZ97], pages 277– 287. ISBN 3-540-63223-9. LCCN QA76.9.D3P16 1997.

Gebhardt:1997:SSB

[Geb97c] Friedrich Gebhardt. Survey on structure-based case retrieval. The Knowledge Engineering Review, 12(1):41–58, 1997. CO-DEN KEREE3. ISSN 0269-8889.

Gebhardt:1995:KEMa

[GGHV95a] Friedrich Gebhardt, Eckehard Groß, Thomas Hemmann, and Hans Voß. Knowledge Engineering mit MoMo [English: Knowledge Engineering with MoMo]. Fabel-Report 29, GMD, Sankt Augustin, Germany, February 1995. 20 pp.

URL ftp://ftp.gmd.de/GMD/
ai-research/Publications/
Fabel/Fabel-29-KE-mit-MoMo.
ps.gz.

Gebhardt:1995:KEMb

[GGHV95b] Friedrich Gebhardt, Eckehard Groß, Thomas Hemmann, and Hans Voß. Knowledge Engineering mit MoMo [English: Knowledge Engineering with MoMo]. KI – Künstliche Intelligenz, 9(1): 22–27, 1995.

Gebhardt:1973:KTD

[GGR73] Friedrich Gebhardt, Hartmann J.

Genrich, and Gernot Richter. Kritik einer Typologie von Daten und Informationssystemen [English: Critique of a typology of data and information systems]. Öffentliche Verwaltung und Datenverarbeitung (ÖVD), 3:524–526, 1973.

Gebhardt:1994:CCCa

[GGV94a] Friedrich Gebhardt, Eckehard Groß, and Hans Voß. Concurrency constraints as control specifications for the MoMo language. Fabel-Report 21, GMD, Sankt Augustin, Germany, May

                                        1994. URL ftp://ftp.gmd.de/
                                        GMD/ai-research/Publications/
                                        Fabel/Fabel-21-MoMo-language.
                                        ps.gz.

Gebhardt:1994:CCCb

[GGV94b] Friedrich Gebhardt, Eckehard Groß, and Hans Voß. Concurrency constraints as control specifications for the MoMo language. In L. Steels, G. Schreiber, and W. Van de Velde, editors, Position papers of the 8th European Knowledge Acquisition Workshop EKAW ‘94. Vrije Universiteit Brussel, Artificial Intelligence Laboratory, Technical Report 94-2, 1994.

Gebhardt:1973:JIE

[GL73] Friedrich Gebhardt and Peter C. Lockemann. Juristisches Informationssystem: Eine Fallstudie für ein Informationssystem großen Umfangs [English: Legal information system: a case study for an information system of a large size]. In Deussen [Deu73], pages 464–472. ISBN 3-540-06127-4.

Gebhardt:1975:MSA

[GLP75] Friedrich Gebhardt, Peter C. Lockemann, and Joachim Poetsch. Möglichkeiten der Suche und Anzeige von Normen und ihren Gliederungseinheiten in Dokumentationssystemen [English: Possibilities for searching and displaying legal norms and their subdivisions in documentation systems]. In Gebhardt [Geb75a], pages 178–193. ISBN 3-8059-0439-8.

Gebhardt:1978:DCD

[GS78] Friedrich Gebhardt and Imant Stellmacher. Design criteria for document retrieval languages. Journal of American Society for Information Science, 29:191– 199, 1978.

Gero:1996:AID

[GS96] John S. Gero and Fay Sudweeks, editors. Artificial Intelligence in Design ’96. Kluwer Academic Publishers Group, Norwell, MA, USA, and Dordrecht, The Netherlands, 1996. ISBN 0-7923-4080-9. LCCN TA174.A7933 1996.

Gutierrez:1991:ASM

[GV91] R. Gutiérrez and M. J. Valderrama, editors. Applied stochastic models and data analysis: proc. 5th International Symposium on ASMDA (Granada, Apr. 1991). World Scientific, Singapore, Philadelphia, 1991. ISBN 981-02-0644-5. LCCN QA402 .I568 1991.

Gebhardt:1997:RCC

[GVGSB97] Friedrich Gebhardt, Angi Voß, Wolfgang Gräther, and Barbara Schmidt-Belz. Reasoning with Complex Cases. Number 393 in The Kluwer Series in Engineering and Computer Science. Kluwer Academic Publishers Group, Norwell, MA, USA, and Dordrecht, The Netherlands, 1997. ISBN 0-7923-9882-3. xi + 250 pp. LCCN Q338.8.R43 1997.

Hossfeld:1976:PRI

[Hoß76] Friedel Hoßfeld, editor. Praxis der Realisierung von Informationssystemen. Number 2 in Applied Computer Science. Carl Hanser, München, Germany, 1976. ISBN 3-446-12236-2. 89– 106 pp.

Komorowski:1997:PDM

[KZ97] Jan Komorowski and Jan Zytkow, editors. Principles of Data Mining and Knowledge Discovery: First European Symposium, PKDD ’97, Trondheim, June 24–27, 1997: Proceedings, volume 1263 of Lecture Notes in Computer Science and Lecture Notes in Artificial Intelligence. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1997. ISBN 3-540-63223-9. LCCN QA76.9.D3P16 1997.

Momirovic:1990:CPC

[MM90] K. Momirovic and V. Mildner, editors. COMPSTAT: proceedings in computational statistics, 9th Symposium (Dubrovnik, Sept. 1990). Physica-Verlag, Heidelberg, Germany, 1990. ISBN 3-7908-0475-4. LCCN QA276.4 .C55 1990.

Muhlbacher:1975:GIF

[Müh75] Jörg Mühlbacher, editor. GI — 5. [i.e. funfte] Jahrestagung: Dortmund, 8.–10. Oktober 1975, volume 34 of Lecture Notes in Computer Science. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1975. CODEN LNCSD9. ISBN 0-387-07410-4. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA267.A1 L43 no.34. English or German.

Muller:1976:BSJ

[Mül76] Bernd S. Müller, editor. Beiträge zur Sprachverarbeitung in juristischen Dokumentationssystemen, volume 7 of Beihefte zu Datenverarbeitung im Recht (DVR). Schweitzer, Berlin, 1976. ISBN 3-8059-0457-6. 202 pp.

Tjoa:1990:DES

[TW90] A Min Tjoa and Roland R. Wagner, editors. Database and expert systems applications: DEXA 90, proceedings of the international conference in Vienna, Austria, August 29–31, 1990. Springer, Wien, 1990. ISBN 3-211-82234-8, 0-387-82234-8. LCCN QA76.9.E96 T551 1990.

Wingert:1978:KFF

[Win78] F. Wingert, editor. Klartextverarbeitung: Frühjahrstagung 1977, Fachbereich Medizinische Informatik der GMDS und Fachausschuß 14 der GI in Gießen, volume 4 of Medizinische Informatik und Statistik. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1978. ISBN 3-540-08634-X.